(6/x) (10/x 4) = 4

Answer by super123(3)

given that
6/x +(10/x+ 4)=4

6(x+ 4)+ 10x
_________ = 4
x(x 4)

(16x+ 24)/(x^2+ 4x)=4
4x^2 +16x=16x+ 24

x^2=6
x= sqrt of 6